HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

RECENT UPDATES TO THE LILLE SPECTROSCOPIC DATABASE

\underline{R. A. MOTIYENKO} , L. MARGULÈS, UMR 8523 - PhLAM - Physique des Lasers Atomes et Molécules, Univ. Lille, CNRS, F-59000 Lille, France

Lille Spectroscopic Database (https://lsd.univ-lille.fr)^{1} hosts the predicted rotational spectra studied by our group in Lille during last 15 years. We will present recent data updates to the database as well as to the application programming interface (API) that allows the integration of the database into third party software. The question of degeneracy factors in case of molecules exhibiting large amplitude motions will be also discussed.

1We thank the Mésocentre de Calcul Scientifique Intensif de l’Université de Lille for hosting the database.

p-number: p228  Submitted on Tue, 20 Jun 2023 21:03:38 +0200